OpenStudy (anonymous):

radical 3x+7=x-1 evaluate

OpenStudy (anonymous):



OpenStudy (anonymous):

3x+7=(x-1)^2 3x+7=x^2-2x+1

OpenStudy (anonymous):

now what do we do?

OpenStudy (anonymous):

5x+7=x2+1

OpenStudy (anonymous):

x^2-5x-6=0

OpenStudy (anonymous):

do i factor?

OpenStudy (anonymous):

(x-6)(x+1)

OpenStudy (anonymous):

x=6,-1 Is it extraneous?

OpenStudy (anonymous):

then we should try answers .

OpenStudy (anonymous):

-1 is not an answer because the answer of square root is positive or zero.

OpenStudy (anonymous):

6 is the answer and it is not extraneous correct?

OpenStudy (anonymous):

yes.(18+7)^(1/2)=6-1 radical25=5